OpenStudy (redohawk):

find the x and y value x = 2 + y -6x - 6y = -12

OpenStudy (redohawk):

@rockinhood if you can show your work this time so i can know how to do it because i have a test soon

OpenStudy (alyssablonde):

This is system of equations correct?

OpenStudy (redohawk):

Correct @AlyssaBlonde

OpenStudy (alyssablonde):

Hmm.. I've never had a problem like this I'm afraid I can't help..

OpenStudy (redohawk):

well, thanks for stopping by :(

OpenStudy (mathstudent55):

Have you learned the substitution method of solving a system of equations?

OpenStudy (redohawk):

ya, it was multiple choice when we learned so i would just plug in the choice they gave until it would work, but there is no multiple choice for this so i need help if you show your work i'd probably be able to understand it

OpenStudy (math&ing001):

Solve the first equation for x, then plug it in the second equation.

OpenStudy (redohawk):

ya, but they both have to match smh

OpenStudy (mathstudent55):

The substitution method works this way: 1. Solve one equation for one variable. 2. Substitute what the variable equals in the other equation. 3. Solve the other equation for one variable. 4. Substitute the known variable in one of the original equations to solve for the other variable.

OpenStudy (mathstudent55):

Read about the substitution method above. Then we'll use it together.

OpenStudy (mathstudent55):

Are you ready?

OpenStudy (redohawk):

i think

OpenStudy (redohawk):

x = 2 + y we can say x = 4 and y =2

OpenStudy (redohawk):

so we solved this one now if you plug in 4 and 2 into the other one it won't match i can't keep guessing and guessing and trying different numbers till one works there has to be a faster way D: HELP!

OpenStudy (mathstudent55):

I don't know. Let's follow the method, one step at a time. 1. We need to solve one equation for one variable. The first equation is already solved for x, so step 1 is already done for us. x = 2 + y

OpenStudy (mathstudent55):

2. Since x = 2 + y, now we substitute x in the second equation with what x is equal to, 2 + y. -6x - 6y = -12 <------ original second equation -6(2 + y) - 6y = -12 <------ 2 + y was plugged in for x

OpenStudy (mathstudent55):

3. Now we have one equation with only one variable, so we can solve for y. -6(2 + y) - 6y = -12 -12 - 6y - 6y = -12 -12 - 12y = -12 -12y = 0 y = 0

OpenStudy (redohawk):

and if y = 0 then x = 2

OpenStudy (mathstudent55):

4. Now that we know that y = 0, we substitute y into either one of the original equations and solve for x. x = 2 + y <---- first original equation x = 2 + 0 <---- y was substituted with 0 x = 2 Solution: x = 2 and y = 0